OpenStudy (anonymous):

How should I evaluate i^-15?

OpenStudy (anonymous):

Ok, it is 1/i^15. U can try to remember that that i^4n =1 so i^4*3 =i^12 =1 So i^15 = i^3 = i* i^2 = -i So -i^-1 is your answer.